in performing a titration using a 0.15 M NaOH as the titrant, the initial reading on a burette is 2.3 ml. The final reading on the burette is 46.2 ml. How many moles of NaOH were used


1
Expert's answer
2021-02-04T07:04:12-0500

∆V=46.2 - 2.3 = 43.9 ml.

Cm=n/V. n=Cm x V = 0.15 x 46.2/1000 = 0,006585 moles.